Suppose that the monetary base is $ 500 billion, the reserve deposit ratio is 0.1 and the currency-deposit ratio is 0.6. Calculate: i. Money Multiplier (2 Marks) ii. Money Supply 


1
Expert's answer
2021-08-12T17:41:48-0400

Solution

a) Money Multiplier

Money Multiplier, Mm "=\frac{1}{reserve. ratio}"

Mm= "=\frac{1}{0.1}= 10"


b) Money Supply

Money supply = "\frac{currency\space ratio +1}{ currency \space ratio + reserve\space ratio}\times M. Base"

"\frac{0.6 +1}{ 0.6 + 0.1}\times 500 = 1,142.857 \space Billion"